# Rigorous results on approach to thermal equilibrium, entanglement, and nonclassicality of an optical quantum field mode scattering from the elements of a non-equilibrium quantum reservoir

### v1

@@ 3.3 The asymptotic system state at weak coupling and approach to equilibrium
$D_{b}(z)={\rm e}^{\varphi(z)},\quad\langle{e^{\varphi(z)}}\rangle_{\sigma}=1+\langle{\varphi(z)}\rangle_{\sigma}+\tfrac{1}{2}\langle{\varphi(z)^{2}}\rangle_{\sigma}+O(|z|^{3}).$
@@
$\chi_{\infty}^{\lambda}(z)=\mathrm{Tr}\big{(}\rho_{\infty}^{\lambda}D_{a}(z)\big{)}=\exp\Big{[}2\lambda^{-1}\langle{\varphi(z)}\rangle_{\sigma}+\tfrac{1}{2}{\rm Var}_{\sigma}\big{(}\varphi(z)\big{)}+O(\lambda)\Big{]},$ (26)
@@
$\chi_{\infty}^{0}(z)=\mathrm{Tr}\big{(}\rho_{\infty}^{0}D_{a}(z)\big{)}:=\lim_{\lambda\rightarrow 0}\mathrm{Tr}\big{(}\rho_{\infty}^{\lambda}D_{a}(z)\big{)}=\exp\Big{[}\tfrac{1}{2}{\rm Var}_{\sigma}\big{(}\varphi(z)\big{)}\Big{]}.$ (27)
@@
${\rm Var}_{\sigma}\big{(}\varphi(z)\big{)}$ $=$ $\text{Cov}_{\sigma}[\varphi(z),\varphi(z)]$ (33)
@@
where $A$ is precisely the covariance matrix of $\sigma$, defined in Eq. (12). Combining this with Eq. (27) we conclude that the state $\rho_{\infty}^{0}$ is the centered Gaussian state having the same covariance matrix $A$ as the single-mode reservoir state $\sigma$.
@@
It then follows again from Eq. (26) that the asymptotic state $\rho^{0}_{\infty,*}$ is the centered Gaussian state with the same covariance matrix as $\sigma$.
@@ 4 Entanglement and (non)classicality of the asymptotic state
$\mathcal{P}^{0}_{\infty}=\lim_{\lambda\to 0}{\rm Tr}(\rho^{\lambda}_{\infty})^{2}=\lim_{\lambda\to 0}{\rm Tr}(\rho_{\infty,*}^{\lambda})^{2}={\rm Tr}\rho_{\textrm{G}}^{2}=\det(V_{\sigma})^{-1/2},$ (37)
@@
${\mathcal{C}}^{2,0}_{\infty}=\lim_{\lambda\to 0}{\mathcal{C}}^{2}(\rho_{\infty}^{\lambda})={\mathcal{C}}^{2}(\rho_{\textrm{G}})\simeq\left(\frac{\mathrm{e}}{2}\right)^{2}{\mathcal{C}}^{2}(\sigma)\exp(-2S_{\infty}^{0}),$
@@
${\mathcal{C}}^{2}(\sigma)=2n+1,\quad{\mathcal{C}}^{0,2}_{\infty}=\frac{1}{2n+1},\quad\mathcal{P}_{\infty}^{0}=\frac{1}{2n+1},\quad S_{\infty}^{0}=g(2n+1).$
@@ A.1 Existence of the asymptotic state
Given a density matrix $\sigma$ and a fixed complex number $z\in\mathbb{C}$, we introduce the function
$\mathbb{R}\ni x\mapsto\chi_{\sigma}(xz)=\langle{D_{b}(xz)}\rangle_{\sigma}\in\mathbb{C}.$
@@
    There is a $\delta>0$ such that for every $z\in\mathbb{C}$ and every $x$ with $|x|<\delta$,
@@
Theorem 1
Suppose that the condition (A1) holds. Then, for all $0<\lambda<\frac{\pi}{2}$, there exists a density matrix $\rho_{\infty}^{\lambda}$ so that for all $z\in\mathbb{C}$,
${\rm Tr}(\rho_{\infty}^{\lambda}D_{a}(z))=\lim_{K\rightarrow\infty}\ \prod_{k=0}^{K-1}\langle{D_{b}(\sin(\lambda)\{\cos(\lambda)\}^{k}z)}\rangle_{\sigma}.$ (40)
Proof. We first show that the limit $K\rightarrow\infty$ in Eq. (16) exists. Recall the abbreviation $s=\sin\lambda$, $c=\cos\lambda$. Let $z\in\mathbb{C}$ be fixed. As $c<1$ there is an integer $k_{0}$ such that $c^{k}<\delta$ for all $k\geq k_{0}$. We use the fundamental theorem of calculus for the function $x\mapsto\ln[\chi_{\sigma}(sxz)]$, to get, for $k\geq k_{0}$,
@@
We have shown so far that the limit as $K\rightarrow\infty$ in Eq. (21) exists. This means that the limit of the characteristic function associated to the density matrix $L^{K}(\rho)$ exists as $K\rightarrow\infty$. Moreover, we have shown that this limit characteristic function, $\chi_{\infty}^{\lambda}$, is continuous in $z$ at the origin $z=0$. It is known [35] that then, $\chi_{\infty}^{\lambda}$ corresponds to a limit density matrix $\rho_{\infty}^{\lambda}$, meaning that there is a density matrix $\rho_{\infty}^{\lambda}$ such that $\chi_{\infty}^{\lambda}(z)={\rm Tr}(\rho_{\infty}^{\lambda}D_{a}(z))$, and that furthermore, $L^{K}(\rho)\rightarrow\rho_{\infty}^{\lambda}$ in trace norm, as $K\rightarrow\infty$. $\square$
@@ A.2 Approach to equilibrium
exists and $\rho_{\infty}^{0}$ is the centered Gaussian state which has the same covariance matrix as the state $\sigma$. Moreover, if the moment $\langle{b}\rangle_{\sigma}$ does not vanish, then $\rho_{\infty}^{\lambda}$ does not have a limit as $\lambda\rightarrow 0$.
@@
Proposition 4
Suppose $\sigma$ satisfies the assumptions (A1) and (A2). For each $z\in\mathbb{C}$ there is a $\lambda_{0}>0$ such that if  $0<\lambda<\lambda_{0}$, then
$\mathrm{Tr}\big{(}\rho_{\infty}^{\lambda}D_{a}(z)\big{)}=\exp\Big{[}2\lambda^{-1}\langle{\varphi(z)}\rangle_{\sigma}+\tfrac{1}{2}{\rm Var}_{\sigma}\big{(}\varphi(z)\big{)}+\lambda t(\lambda,z)\Big{]},$ (47)
@@
By condition (A2), the map $z\mapsto{\rm Var}_{\sigma}(\varphi(z))$ is continuous at the origin. Eq. (48) means that the characteristic function of $\rho_{\infty}^{\lambda}$ has a limit as $\lambda\rightarrow 0$, and this limit is continuous in $z$ at the origin $z=0$. It follows from [35] (“SWOT convergence Lemma”) that $\rho_{\infty}^{\lambda}$ converges in trace norm to some density matrix we denote $\rho_{\infty}^{0}$, as $\lambda\rightarrow 0$, and that moreover, the characteristic function of $\rho_{\infty}^{0}$ is the limit characteristic function of the $\rho_{\infty}^{\lambda}$. In other words, for all $z\in\mathbb{C}$,
@@
This shows that $\rho_{\infty}^{0}$ is Gaussian. By proceeding as in Eq. (33)-(34), we identify $\rho_{\infty}^{0}$ as the centered Gaussian having the same covariance matrix $A$ as $\sigma$. This completes the the proof of Theorem 2, modulo a proof of Proposition 4, which we give now.
@@
$\mathrm{Tr}\big{(}\rho_{\infty}^{\lambda}D_{a}(z)\big{)}=\prod_{k\geq 0}\langle{{\rm e}^{\varphi(sc^{k}z)}}\rangle_{\sigma}=\exp\big{[}\sum_{k\geq 0}\ln\langle{{\rm e}^{\varphi(\epsilon_{k}z)}}\rangle_{\sigma}\big{]},$
@@
$\ln\langle{{\rm e}^{\varphi(\epsilon_{k}z)}}\rangle_{\sigma}$ $=$ $\ln\big{[}1+\epsilon_{k}\langle{\varphi(z)}\rangle_{\sigma}+\tfrac{\epsilon_{k}^{2}}{2}\langle{\varphi(z)^{2}}\rangle_{\sigma}+R_{\sigma}(\epsilon_{k}z)\big{]}$ (50)
@@
$\Big{|}\xi_{k}-\tfrac{1}{2}\xi_{k}^{2}-\big{\{}\epsilon_{k}\langle{\varphi(z)}\rangle_{\sigma}+\tfrac{1}{2}\epsilon_{k}^{2}{\rm Var}_{\sigma}(\varphi(z))\big{\}}\Big{|}\leq s^{3}c^{3k}C(z)$ (55)
@@
$\Big{|}\ln\langle{{\rm e}^{\varphi(\epsilon_{k}z)}}\rangle_{\sigma}-\big{\{}\epsilon_{k}\langle{\varphi(z)}\rangle_{\sigma}+\tfrac{1}{2}\epsilon_{k}^{2}{\rm Var}_{\sigma}(\varphi(z))\big{\}}\Big{|}\leq s^{3}c^{3k}C(z)$ (56)
@@
$\Big{|}\sum_{k=0}^{K}\Big{[}\ln\langle{{\rm e}^{\varphi(sc^{k}z)}}\rangle_{\sigma}-\big{\{}sc^{k}\langle{\varphi(z)}\rangle_{\sigma}+\tfrac{1}{2}s^{2}c^{2k}{\rm Var}_{\sigma}(\varphi(z))\big{\}}\Big{]}\Big{|}\leq s^{3}C(z)\sum_{k=0}^{K}c^{3k}.$ (57)
@@
$\sum_{k\geq 0}\ln\langle{{\rm e}^{\varphi(sc^{k}z)}}\rangle_{\sigma}=\frac{s}{1-c}\langle{\varphi(z)}\rangle_{\sigma}+\tfrac{1}{2}{\rm Var}_{\sigma}(\varphi(z))+t(s,z)$

### v2

@@ 3.3 The asymptotic system state at weak coupling and approach to equilibrium
$D_{b}(z)$ $=$ ${\rm e}^{\varphi(z)},$
$\langle{e^{\varphi(z)}}\rangle_{\sigma}$ $=$ $1+\langle{\varphi(z)}\rangle_{\sigma}+\tfrac{1}{2}\langle{\varphi(z)^{2}}\rangle_{\sigma}+O(|z|^{3}).$
@@
$\chi_{\infty}^{\lambda}(z)=\mathrm{Tr}\big{(}\rho_{\infty}^{\lambda}D_{a}(z)\big{)}$
$=$ $\exp\Big{[}2\lambda^{-1}\langle{\varphi(z)}\rangle_{\sigma}+\tfrac{1}{2}{\rm Var}_{\sigma}\big{(}\varphi(z)\big{)}+O(\lambda)\Big{]},$
@@
$\chi_{\infty}^{0}(z)$ $=$ $\mathrm{Tr}\big{(}\rho_{\infty}^{0}D_{a}(z)\big{)}$
::=:= $\lim_{\lambda\rightarrow 0}\mathrm{Tr}\big{(}\rho_{\infty}^{\lambda}D_{a}(z)\big{)}$
$=$ $\exp\Big{[}\tfrac{1}{2}{\rm Var}_{\sigma}\big{(}\varphi(z)\big{)}\Big{]}.$
@@
${\rm Var}_{\sigma}\big{(}\varphi(z)\big{)}=\text{Cov}_{\sigma}[\varphi(z),\varphi(z)]$
@@
where $A$ is precisely the covariance matrix of $\sigma$, defined in Eq. (12). Combining this with Eq. (3.3) we conclude that the state $\rho_{\infty}^{0}$ is the centered Gaussian state having the same covariance matrix $A$ as the single-mode reservoir state $\sigma$.
@@
It then follows again from Eq. (3.3) that the asymptotic state $\rho^{0}_{\infty,*}$ is the centered Gaussian state with the same covariance matrix as $\sigma$.
@@ 4 Entanglement and (non)classicality of the asymptotic state
$\mathcal{P}^{0}_{\infty}$ $=$ $\lim_{\lambda\to 0}{\rm Tr}(\rho^{\lambda}_{\infty})^{2}=\lim_{\lambda\to 0}{\rm Tr}(\rho_{\infty,*}^{\lambda})^{2}$ (33)
$=$ ${\rm Tr}\rho_{\textrm{G}}^{2}=\det(V_{\sigma})^{-1/2},$
@@
${\mathcal{C}}^{2,0}_{\infty}$ $=$ $\lim_{\lambda\to 0}{\mathcal{C}}^{2}(\rho_{\infty}^{\lambda})={\mathcal{C}}^{2}(\rho_{\textrm{G}})$
$\simeq$ $\left(\frac{\mathrm{e}}{2}\right)^{2}{\mathcal{C}}^{2}(\sigma)\exp(-2S_{\infty}^{0}),$
@@
${\mathcal{C}}^{2}(\sigma)=2n+1,\quad{\mathcal{C}}^{0,2}_{\infty}=\frac{1}{2n+1},$
$\mathcal{P}_{\infty}^{0}=\frac{1}{2n+1},\quad S_{\infty}^{0}=g(2n+1).$
@@ A.1 Existence of the asymptotic state
Given a density matrix $\sigma$ and a fixed complex number $z\in\mathbf{C}$, we introduce the function
$\mathbf{R}\ni x\mapsto\chi_{\sigma}(xz)=\langle{D_{b}(xz)}\rangle_{\sigma}\in\mathbf{C}.$
@@
    There is a $\delta>0$ such that for every $z\in\mathbf{C}$ and every $x$ with $|x|<\delta$,
@@
Theorem 1.
Suppose that the condition (A1) holds. Then, for all $20<\lambda<\pi/$ 2, there exists a density matrix $\rho_{\infty}^{\lambda}$ so that for all $z\in\mathbf{C}$,
${\rm Tr}(\rho_{\infty}^{\lambda}D_{a}(z))=$ (36)
$\lim_{K\rightarrow\infty}\ \prod_{k=0}^{K-1}\langle{D_{b}(\sin(\lambda)\{\cos(\lambda)\}^{k}z)}\rangle_{\sigma}.$
Proof. We first show that the limit $K\rightarrow\infty$ in Eq. (17) exists. Recall the abbreviation $s=\sin\lambda$, $c=\cos\lambda$. Let $z\in\mathbf{C}$ be fixed. As $c<1$ there is an integer $k_{0}$ such that $c^{k}<\delta$ for all $k\geq k_{0}$. We use the fundamental theorem of calculus for the function $x\mapsto\ln[\chi_{\sigma}(sxz)]$, to get, for $k\geq k_{0}$,
[…]
@@
We have shown so far that the limit as $K\rightarrow\infty$ in Eq. (22) exists. This means that the limit of the characteristic function associated to the density matrix $L^{K}(\rho)$ exists as $K\rightarrow\infty$. Moreover, we have shown that this limit characteristic function, $\chi_{\infty}^{\lambda}$, is continuous in $z$ at the origin $z=0$. It is known [41] that then, $\chi_{\infty}^{\lambda}$ corresponds to a limit density matrix $\rho_{\infty}^{\lambda}$, meaning that there is a density matrix $\rho_{\infty}^{\lambda}$ such that $\chi_{\infty}^{\lambda}(z)={\rm Tr}(\rho_{\infty}^{\lambda}D_{a}(z))$, and that furthermore, $L^{K}(\rho)\rightarrow\rho_{\infty}^{\lambda}$ in trace norm, as $K\rightarrow\infty$. ∎
@@ A.2 Approach to equilibrium
exists and $\rho_{\infty}^{0}$ is the centered Gaussian state which has the same covariance matrix as the state $\sigma$.
Moreover, if the moment $\langle{b}\rangle_{\sigma}$ does not vanish, then $\rho_{\infty}^{\lambda}$ does not have a limit as $\lambda\rightarrow 0$.
@@
Proposition 4.
Suppose $\sigma$ satisfies the assumptions (A1) and (A2). For each $z\in\mathbf{C}$ there is a $\lambda_{0}>0$ such that if  $0<\lambda<\lambda_{0}$, then
$\mathrm{Tr}\big{(}\rho_{\infty}^{\lambda}D_{a}(z)\big{)}$
$=$ $\exp\Big{[}2\lambda^{-1}\langle{\varphi(z)}\rangle_{\sigma}+\tfrac{1}{2}{\rm Var}_{\sigma}\big{(}\varphi(z)\big{)}+\lambda t(\lambda,z)\Big{]},$
@@
By condition (A2), the map $z\mapsto{\rm Var}_{\sigma}(\varphi(z))$ is continuous at the origin. Eq. (44) means that the characteristic function of $\rho_{\infty}^{\lambda}$ has a limit as $\lambda\rightarrow 0$, and this limit is continuous in $z$ at the origin $z=0$. It follows from [41] (“SWOT convergence Lemma”) that $\rho_{\infty}^{\lambda}$ converges in trace norm to some density matrix we denote $\rho_{\infty}^{0}$, as $\lambda\rightarrow 0$, and that moreover, the characteristic function of $\rho_{\infty}^{0}$ is the limit characteristic function of the $\rho_{\infty}^{\lambda}$. In other words, for all $z\in\mathbf{C}$,
@@
This shows that $\rho_{\infty}^{0}$ is Gaussian. By proceeding as in Eq. (3.3)-(30), we identify $\rho_{\infty}^{0}$ as the centered Gaussian having the same covariance matrix $A$ as $\sigma$. This completes the the proof of Theorem 2, modulo a proof of Proposition 4, which we give now.
@@
$\mathrm{Tr}\big{(}\rho_{\infty}^{\lambda}D_{a}(z)\big{)}$ $=$ $\prod_{k\geq 0}\langle{{\rm e}^{\varphi(sc^{k}z)}}\rangle_{\sigma}$
$=$ $\exp\big{[}\sum_{k\geq 0}\ln\langle{{\rm e}^{\varphi(\epsilon_{k}z)}}\rangle_{\sigma}\big{]},$
@@
$\ln\langle{{\rm e}^{\varphi(\epsilon_{k}z)}}\rangle_{\sigma}$ $=$ $\ln\big{[}1+\epsilon_{k}\langle{\varphi(z)}\rangle_{\sigma}+\tfrac{\epsilon_{k}^{2}}{2}\langle{\varphi(z)^{2}}\rangle_{\sigma}$ (46)
$\quad+R_{\sigma}(\epsilon_{k}z)\big{]}$
@@
$\Big{|}\xi_{k}-\tfrac{1}{2}\xi_{k}^{2}-\big{\{}\epsilon_{k}\langle{\varphi(z)}\rangle_{\sigma}+\tfrac{1}{2}\epsilon_{k}^{2}{\rm Var}_{\sigma}(\varphi(z))\big{\}}\Big{|}\\ \leq s^{3}c^{3k}C(z)$ (51)
@@
$\Big{|}\ln\langle{{\rm e}^{\varphi(\epsilon_{k}z)}}\rangle_{\sigma}-\big{\{}\epsilon_{k}\langle{\varphi(z)}\rangle_{\sigma}+\tfrac{1}{2}\epsilon_{k}^{2}{\rm Var}_{\sigma}(\varphi(z))\big{\}}\Big{|}\\ \leq s^{3}c^{3k}C(z)$ (52)
@@
$\Big{|}\sum_{k=0}^{K}\Big{[}\ln\langle{{\rm e}^{\varphi(sc^{k}z)}}\rangle_{\sigma}-\big{\{}sc^{k}\langle{\varphi(z)}\rangle_{\sigma}\\ +\tfrac{1}{2}s^{2}c^{2k}{\rm Var}_{\sigma}(\varphi(z))\big{\}}\Big{]}\Big{|}\leq s^{3}C(z)\sum_{k=0}^{K}c^{3k}.$ (53)
@@
$\sum_{k\geq 0}\ln\langle{{\rm e}^{\varphi(sc^{k}z)}}\rangle_{\sigma}=\\ \frac{s}{1-c}\langle{\varphi(z)}\rangle_{\sigma}+\tfrac{1}{2}{\rm Var}_{\sigma}(\varphi(z))+t(s,z)$